Thinfilms morning report 04/08/03

Jim Haydon jhaydon at snf.stanford.edu
Tue Apr 8 10:18:19 PDT 2003




Balzers:   Blazers will be taken down some time this week to fix a water
leak in the RF feed through an shield cleaning . It will be down for at
least a week
Gryphon:    ok
Implanter:    ok
Innotec:     ok
Metalica:    ok
STS Dep:    The RF is working now. It seems to have a flakey switch when
selecting between high and low frequencies. It probably will show up as
a green RF on LED on the low frequency power supply. This light should
be on only when it is being asked.
wbgen- CTB:    ok
wbgen-LHP:    ok
wbgen-RHP:    ok
wbMETAL:    ok
wbNITRIDE:    ok
wbNONMETAL:    ok
wbDIFF:    ok
wbSILICIDE:    ok
wbGAASCL:    ok
wbMISCRES:    ok
wbOPTRES:    ok
DI WATER:    ok
NITROGEN:    ok





More information about the thinfilms mailing list